OpenStudy (anonymous):

4x + 5y = 7 3x – 2y = –12

OpenStudy (jhannybean):

Process of elimination. Eliminate one variable and then use it to find the other.

OpenStudy (anonymous):

Thanks

OpenStudy (jhannybean):

4x + 5y = 7 3x – 2y = –12 Let's eliminate out x's first. Common multiple is 12, so multiply the first equation by -3 and the second equation by 4 -3(4x +5y = 7) 4(3x-2y = 12) What do you get?

OpenStudy (anonymous):

-2,3

OpenStudy (anonymous):

you can eliminate y instead of x too right?